Find the vertex and focus of the parabola : y^2 + 4y + 8x -12=0

Answer by ewatrrr(24785)

y^2 + 4y + 8x -12=0
(y + 2)^2 - 4 + 8x -12 = 0
(y+2)^2 = -8x + 16 V(
(y+2)^2 = -8(x-2) V(2, -2) Opening left along y = -2
4p = -8, p = -2, F(0,-2)